OpenStudy (anonymous):

Use the zero-product property to solve the equation: (x + 4)(x - 2) = 0

OpenStudy (anonymous):

I don't understand how to use the zero-product property.

OpenStudy (anonymous):

let (x+4) = 0 and (x-2) = 0 now solve each equation individually

OpenStudy (tkhunny):

The Zero-Product property is in interesting little fellow. What do we know about A and B if we know that A*B = 5? We know almost nothing. What do we know about A and B if we know that A*B = 0? We know that either A = 0 or B = 0. That's WAY more than anything that isn't zero.